OpenStudy (anonymous):

I've asked this before but I wasn't able to figure it out. I thought I had it so I closed the question. Maybe a little more help? x^2+2x+9=0

OpenStudy (anonymous):

x^2+2x+1+8=0 (x+1)^2+8=0 (x+1)^2=-8 x+1=sqrt(-8)=2sqrt2*i x=2sqrt2*i-1

OpenStudy (mathstudent55):

Since the left side can't be factored, you can complete the square or use the quadratic formula.

OpenStudy (anonymous):

Oh, wow. Thanks! I kept getting \[x=-1\pm \sqrt{-18}\] and I know that is not right

OpenStudy (anonymous):

My answer choices are a. \[x=-2\pm4i \sqrt{2}\] b. \[x=-2\pm2i \sqrt{2}\] c. \[x=-1\pm4i \sqrt{2}\] d. \[x=-1\pm2i \sqrt{2}\]

OpenStudy (anonymous):

I think it may be C...?

OpenStudy (anonymous):

i said 2sqrt2*i-1. so answer is d :)

OpenStudy (anonymous):

Oh! I'm sorry, I looked at it wrong :) I have it worked out wrong on my paper so I got confuzzled ;P

OpenStudy (mathstudent55):